An object is released at an initial velocity of zero in the earth’s gravitational field. what is its velocity after 3 seconds ?
Answer #1

vf = vi + at

or, vf = 0 + gt , put t = 3 s

or, vf = 29.4 m/s

***********************************************************************************************
Check the answer and let me know immediately if you find something wrong or missing... I will rectify the mistakes asap if any